OpenStudy (anonymous):

a1+a2+a3…+an is an arithmetic series, what is the sum?

OpenStudy (shubhamsrg):

The sum is just what you wrote! ^_^

OpenStudy (anonymous):

all we have to do find the sum of first n of AP series with first term as a1 and common difference d ( which u can calculate by subtracting next term with its previous term ex: d = a2 - a1 ) now sum is \[\sum = \frac{ n }{ 2 }(2a+(n-1)d)\]

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

Its oKay